A) What is X if X equals the value of investment A plus the value of investment B? Investment A is expected to pay 16,000 dollars in 1 year(s) from today and has an expected return of 4.23 percent per year. Investment B is expected to pay 14,300 dollars in 2 year(s) from today and has an expected return of 5.12 percent per year.

B) Sasha owns two investments, A and B, that have a combined total value of 40,500 dollars. Investment A is expected to pay 25,400 dollars in 7 year(s) from today and has an expected return of 8.39 percent per year. Investment B is expected to pay X in 6 years from today and has an expected return of 5.43 percent per year. What is X, the cash flow expected from investment B in 6 years from today?

C) Sasha owns two investments, A and B, that have a combined total value of 59,000 dollars. Investment A is expected to pay 26,300 dollars in 5 year(s) from today and has an expected return of 5.69 percent per year. Investment B is expected to pay 63,851 in 6 years from today and has an expected return of R per year. What is R, the expected annual return for investment B? Answer as a rate in decimal format so that 12.34% would be entered as .1234 and 0.98% would be entered as .0098.
Answer #1

A). X is present value of both the investment

PV = FV/(1+r)^t

X = PV of investment A + PV of investment B

X = 16000/(1+0.0423) + 14300/(1+0.0512)^2 => X = $28269.46

B). Combined investment = $40500

Investment A is expected to pay 25,400 dollars in 7 year(s) from today and has an expected return of 8.39 percent per year. Investment B is expected to pay X in 6 years from today and has an expected return of 5.43 percent per year.

So, 40500 = 25400/(1+0.0839)^7 + X/(1+0.0543)^6

=> X = $35774.21

C). Combined investment = $59000

Investment A is expected to pay 26,300 dollars in 5 year(s) from today and has an expected return of 5.69 percent per year. Investment B is expected to pay 63,851 in 6 years from today and has an expected return of R per year.

So, 59000 = 26300/(1+0.0569)^5 + 63851/(1+R)^6

=> R = 8.54% or 0.0854